I just want to make sure I got this one right.
2x^2-7x=4
2x^2-7x-4=0
(2x+4)(7x-4)=0
2x+4=0 7x-4=0
2x=-4
x=-1/2 x=4
Can any body tell me if I got it right.

Answer by Alan3354(69443)

I just want to make sure I got this one right.
2x^2-7x=4
2x^2-7x-4=0
(2x+4)(7x-4)=0 **** If you multiply this out, you get 14x^2+20x-16
--------------
(2x+1)*(x-4) = 0
x = -1/2, x = 4
------------------
------------------
2x+4=0 7x-4=0
2x=-4 --> x = -2 How did you get -1/2 ?
And 7x-4 = 0 --> x = 4/7
x=-1/2 x=4
Can any body tell me if I got it right.
You got the right answers, I suspect they were in the book.